6. A force of 20 lb stretches a spring from a natural length of 7 inches to a length of 12 inches. How much work was performed in stretching the spring to this length?

1
Expert's answer
2022-05-30T16:56:35-0400
"F=k\Delta x"

"k=\dfrac{F}{x_2-x_1}"

"=\dfrac{20\ lb}{12\ in-7\ in}=4\ lb/in"

"W=\dfrac{k(\Delta x)^2}{2}"

"=\dfrac{4\ lb/in(12\ in-7\ in)^2}{2}=50\ lb\cdot in"